OpenStudy (anonymous):

A coin is tossed 3 times in a row. What is the probability that heads appears on only the last toss.

OpenStudy (anonymous):

HHH , HHT , HTH , THH , HTT , THT , TTH , TTT I thought it was 3/8 but that was wrong.

OpenStudy (kropot72):

First toss: P(tail) = 1/2 Second toss: P(tail) = 1/2 Last toss: P(head) = 1/2 The result of each toss is independent of the other results. Therefore the required probability is given by: \[P(head\ only\ on\ last\ toss)=\frac{1}{2}\times\frac{1}{2}\times\frac{1}{2}=you\ can\ calculate\]

OpenStudy (anonymous):

its 1/8 ☻☺